OpenStudy (johndp123456):

pls help 13+w/7=-18 will fan and medal

OpenStudy (johndp123456):

@Awolflover1

Awolflover1 (awolflover1):

Hii (:

OpenStudy (johndp123456):

need help with one more

Awolflover1 (awolflover1):

\[13+\frac{ w }{ 7 }=-18\]

Awolflover1 (awolflover1):

so Simplify \[\frac{ 1 }{ 7 }w+13=-18\] then subtract 13 on both sides \[\frac{ 1 }{ 7 }+13-13=-18-13\]

OpenStudy (johndp123456):

so you would have 1/7=-31

Awolflover1 (awolflover1):

yep :) but we have to get w

Awolflover1 (awolflover1):

Multiply both sides by 7 \[7\times(\frac{ 1 }{ 7 }w)=(7)\times(-31)\]

OpenStudy (johndp123456):

1w=-217

Awolflover1 (awolflover1):

yes there we go!!

OpenStudy (johndp123456):

thank you so much ;)

OpenStudy (johndp123456):

you made my day better

Awolflover1 (awolflover1):

haha you're welcome! tag me if you need more help i will try my best :)